The ice point and steam point on a thermometer are found to be 96.0mm apart. What temperature is recorded when the length of mercury thread is 33.6mm above the ice point mark
1
Expert's answer
2020-04-29T09:48:25-0400

As per the question,

The ice point on the thermometer =96.0mm

recorded length of the mercury =33.6mm

Let the required temperature is x,


"\dfrac{x-0}{100-0}=\dfrac{33.6-0}{96-0}"

"\Rightarrow x=\dfrac{33.6\times 100}{96}^\circ C"

"x=35^\circ C"